OpenStudy (anonymous):

Please explain step by step The function h = -t2 + 95 models the path of a ball thrown by a boy where h represents height, in feet, and t represents the time, in seconds, that the ball is in the air. Assuming the boy lives at sea level where h = 0 ft, which is a likely place the boy could have been standing when he threw this ball? A. his backyard B. an underground cave C. a ladder D. a bridge

OpenStudy (anonymous):

At t=0, what is the height of the ball?

OpenStudy (anonymous):

um... 0? e.o

OpenStudy (anonymous):

\[\large h(0)=-0^2 +95 = ?\]

OpenStudy (anonymous):

The equation: h = -t^2 + 95 Describes the height of the ball as a function of time (t). When t is zero, we have: h = -(0)^2 + 95 h = 95 feet That is, before he threw the ball, the height is at 95 feet. That is about the height of a bridge. Answer: d. a bridge

OpenStudy (anonymous):

Oooh that makes sense... thank you both.

OpenStudy (anonymous):

You're welcome!! Xo!